OpenStudy (anonymous):

John deposited $2860 in a bank that pays 9% interest, compounded monthly. Find the amount he will have at the end of 3 years.

OpenStudy (jdoe0001):

\(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ A=\textit{current amount}\\ P=\textit{original amount deposited}\to \$2,860\\ r=rate\to 9\%\to \frac{9}{100}\to 0.09\\ n=\textit{compounding cycle, monthly, means 12 times per year}\to 12\\ t=years\to 3 \\ \quad \\ A=2,860\left(1+\frac{0.09}{12}\right)^{12\cdot 3}\)